# Neighborhood Feature Pooling for Remote Sensing Image Classification

### v1

@@ Abstract
In this work, we propose neighborhood feature pooling (NFP) as a novel texture feature extraction method for remote sensing image classification. The NFP layer captures relationships between neighboring inputs and efficiently aggregates local similarities across feature dimensions. This new approach is implemented using convolutional layers and can be seamlessly integrated into any network. Results comparing the baseline models and the NFP method indicate the potential of this new approach for classification. NFP consistently improves performance across diverse datasets and architectures while maintaining minimal parameter overhead. The code for this work is publicly available. [^1] https://github.com/Advanced-Vision-and-Learning-Lab/Neighbour_Feature_Pooling
@@ 1 Introduction
However, standard CNN architectures can inadvertently lose these cues when using traditional pooling operations (e.g., average and max pooling). These layers are effective in summarizing information at reduced spatial dimensions, yet typically fail to encode important structural details [33]. This limitation stems from the lack of explicit weighting on neighboring relationships within the feature space. To address this gap, neighborhood feature pooling (NFP) is introduced in this work. This is a novel layer that measures how similar each neighbor is to a center pixel and/or feature vector. NFP can work alongside the usual global average pooling (GAP), adding a texture-aware branch for improved feature representation.
@@ 2 Related Work
(b) Detailed illustration of NFP with $r=1$. An input patch from the UC Merced dataset (“buildings” class) is processed with eight directional difference kernels, which measure relative feature differences between the center pixel and its neighbors (scaled dot product Similarity). For visualization, the outputs are normalized to the range [0,1] and computed with a dilation factor of 15, yielding NFP maps that form a multi-channel similarity representation of local texture relationships, later aggregated into the NFP feature vector.
Figure 1: NFP illustration. (a) Overview of the Neighborhood Feature Pooling framework. (b) Example with $r=1$, showing the input patch, directional kernels, and the resulting similarity maps that encode local texture relationships. Together these demonstrate how NFP captures local structural information that is later integrated with global features (Fig. 2).
@@
Texture descriptors in images may be statistical, summarizing pixel-value distributions, or structural, modeling relationships among neighboring pixels [19]. Two recent structural approaches exploit similarity maps to enrich convolutional features. The Local Similarity Pattern (LSP) layer appends cosine-similarity maps between each pixel and its eight-neighbor window, supplying additional structural cues for stereo-matching networks [14]. In multi-sensor anomalous change detection, a Neighborhood Similarity Feature Space (NSFS) concatenates dot-product similarity responses as extra channels, boosting texture discrimination without resampling [9]. While these methods demonstrate the benefit of neighbor-wise similarity, they employ fixed similarity metrics and rely on pooling operations that can reduce fine-grained texture patterns.
These advances highlight the importance of neighborhood similarity for texture encoding, but current methods remain constrained by handcrafted descriptors, fixed similarity choices, or stochastic sampling. This motivates our proposed Neighborhood Feature Pooling (NFP), which integrates learnable similarity maps directly into the backbone for end-to-end optimization with minimal parameter overhead.
@@ 3 Method
Figure 2: Full architecture illustration of the proposed model using MobileNetV3 [10] with NFP. Each block represents a stage in the feature extraction pipeline. After feature maps are extracted from the input, the features are aggregated through two branches: global average pooling (GAP) and NFP. The NFP branch first extract the similarity maps then the similarity values are aggregated through GAP. The average NFP features are then upsampled to the same dimension using a $1\times 1$ convolution. The final step is for the GAP and NFP feature vectors to be multiplied before being passed into the output classification layer.
@@ 3.1 Neighborhood Similarity Computation
where $d(\cdot,\cdot)$ is a similarity function such as cosine or dot product. This operation yields a set of $(2r+1)^{2}-1$ scalar values per center pixel, producing a multi-channel similarity map that encodes local texture structure, as illustrated in Fig. 1.
Similarity measures such as cosine or Pearson correlation can yield negative values, which are retained to represent dissimilar or anti-correlated relationships between neighbors. For distance-based metrics such as $L_{p}$ norm or RMSE, the computed distances are negated so that higher similarity consistently corresponds to larger values. This convention provides a unified interpretation across metrics and ensures that all similarity and distance functions are handled within a common framework.
In our experiments we fix $r=1$, which results in 8 similarity values per feature channel. Larger neighborhoods could be used to capture broader spatial context at the expense of reduced boundary precision. The formulation in Eq. 1 thus provides an explicit mechanism for encoding relative feature-space relationships across local neighborhoods, preserving structural cues often discarded by global pooling.
@@ 3.2 Similarity Function Options
The similarity function $d(\cdot,\cdot)$ in Eq. 1 can be drawn from a wide set of metrics, each emphasizing different aspects of feature space geometry. The choice of metric can influence NFP by defining how local neighborhood relationships are quantified. We follow the taxonomy of the first three categories defined by Deborah et al. [5] as the feature vectors are similar to hyperspectral signatures used in remote sensing. We also introduced additional similarity metrics inspired by recent deep learning methods:
@@
Backbone Method PlantVillage RESISC45 UC Merced GTOS-Mobile EuroSAT
ResNet18 GAP 99.68 ± 0.07 (11.20) 91.28 ± 0.04 (11.20) 97.62 ± 0.19 (11.19) 72.01 ± 0.01 (11.19) 98.27 ± 0.19 (11.21)
Lacunarity 99.73 ± 0.04 (11.20) 93.10 ± 0.01 (11.20) 97.30 ± 0.11 (11.19) 73.12 ± 3.21 (11.19) 98.23 ± 0.33 (11.21)
Fractal 99.68 ± 0.03 (11.45) 92.29 ± 0.16 (11.47) 97.62 ± 0.34 (11.45) 69.43 ± 2.27 (11.46) 98.49 ± 0.20 (11.47)
RADAM 99.61 ± 0.10 (11.20) 92.08 ± 0.15 (11.20) 97.46 ± 0.22 (11.20) 69.37 ± 2.68 (11.20) 98.40 ± 0.23 (11.21)
DeepTEN 99.25 ± 0.05 (11.47) 93.19 ± 0.05 (11.96) 97.14 ± 0.58 (11.57) 71.66 ± 0.07 (11.73) 98.50 ± 0.28 (11.42)
NFP (Ours) 99.76 ± 0.01 (11.23) 93.22 ± 0.25 (11.23) 98.86 ± 0.58 (11.23) 74.52 ± 0.02 (11.23) 98.52 ± 0.20 (11.26)
MobileNetV3 GAP 99.83 ± 0.04 (4.22) 94.48 ± 0.37 (4.23) 97.57 ± 0.57 (4.22) 76.27 ± 1.49 (4.23) 98.23 ± 0.13 (4.21)
Lacunarity 99.82 ± 0.07 (4.22) 92.93 ± 0.41 (4.25) 97.94 ± 0.59 (4.23) 73.81 ± 0.26 (4.23) 98.52 ± 0.32 (4.21)
Fractal 99.76 ± 0.11 (5.14) 93.55 ± 0.19 (5.17) 97.94 ± 0.22 (5.15) 75.46 ± 0.12 (5.15) 98.35 ± 0.15 (5.14)
RADAM 99.45 ± 0.05 (4.22) 85.59 ± 2.61 (4.22) 97.78 ± 1.81 (4.22) 75.89 ± 3.58 (4.22) 98.27 ± 0.05 (4.21)
DeepTEN 99.73 ± 0.08 (4.76) 94.95 ± 0.09 (5.68) 98.49 ± 0.31 (4.93) 77.51 ± 0.23 (5.25) 98.35 ± 0.23 (4.60)
NFP (Ours) 99.86 ± 0.02 (4.31) 94.80 ± 0.21 (4.31) 98.49 ± 0.28 (4.31) 76.49 ± 1.17 (4.31) 98.53 ± 0.12 (4.30)
ViT-Tiny GAP 99.62 ± 0.06 (5.53) 93.06 ± 0.49 (5.54) 97.78 ± 0.11 (5.53) 68.64 ± 3.75 (5.53) 97.96 ± 0.14 (5.53)
Lacunarity 99.51 ± 0.04 (5.53) 93.42 ± 0.03 (5.54) 97.86 ± 0.19 (5.53) 74.71 ± 2.67 (5.53) 89.20 ± 0.36 (6.02)
Fractal 99.62 ± 0.03 (5.57) 92.57 ± 0.61 (5.58) 97.86 ± 0.39 (5.57) 67.33 ± 0.78 (5.57) 92.47 ± 0.35 (6.05)
RADAM 99.19 ± 0.29 (5.53) 91.75 ± 0.23 (5.53) 96.67 ± 0.89 (5.53) 67.68 ± 1.74 (5.53) 91.83 ± 1.21 (6.02)
DeepTEN 99.25 ± 0.14 (5.64) 93.17 ± 0.09 (5.82) 97.62 ± 0.34 (5.67) 69.84 ± 2.66 (5.73) 88.42 ± 1.72 (6.09)
NFP (Ours) 99.57 ± 0.07 (5.55) 92.79 ± 0.17 (5.55) 97.38 ± 0.34 (5.55) 75.47 ± 0.05 (5.55) 94.51 ± 0.88 (6.03)
Table 1: Classification accuracy (%) and model size (in millions of parameters) across five datasets using various texture pooling methods. Each cell shows Accuracy $\pm$ Standard Deviation, followed by parameter count in parentheses. For each backbone architecture, the best-performing method on each dataset is highlighted in bold.
@@ 3.3 Model Architecture Integration
The affinity stack is spatially averaged and projected to a vector in $\mathbb{R}^{C^{\prime}}$ with $C^{\prime}=C$. A $1\times 1$ convolution is applied only if $C^{\prime}\neq C$. The resulting vector is fused with the backbone’s GAP representation via element-wise multiplication. Because both streams are pooled to vectors, the difference between $(H,W)$ and $(H^{\prime},W^{\prime})$ does not affect the fusion. This design preserves pretrained backbone weights and introduces only a lightweight projection layer as additional learnable overhead, making NFP broadly applicable across architectures such as ResNet, MobileNetV3, and ViT-Tiny, as detailed in Sec. 4.
4 Experiment
@@ 4.1 Experimental setup
Five public remote sensing datasets are used for evaluation: PlantVillage [11], UC Merced Land Use [30], RESISC45 [3], Ground Terrain in Outdoor Scenes (GTOS) Mobile [29], and EuroSAT. The first four datasets were RGB while EuroSAT was multi-spectral with 13 bands. PlantVillage is used with an 80/10/10 split, GTOS-Mobile follows the official split (84.5/9.4/6.1), and UC Merced, RESISC45, and EuroSAT follow the standard train/validation/test protocols described by Neumann et al. [17]. To ensure fairness across baselines, all models are trained using identical data splits and preprocessing pipelines.
@@ 4.2 Backbone Networks
The NFP layer is evaluated across three backbone architectures: ViT-Tiny [6], MobileNetV3-Large [10], and ResNet-18 [27]. All models are initialized with pre-trained weights from the PyTorch Image Models (timm) library [26] and fine-tuned on each dataset. The NFP module is integrated into each backbone according to the scheme described in Sec. 3.3. In all cases, NFP is placed after the final feature extraction stage, followed by a linear classification head. A $1{\times}1$ convolution is applied after the NFP to align the number of feature maps from NFP to the average pooled feature maps. This design ensures compatibility with the classifier while maintaining low parameter overhead. Classification is performed by a linear head after multiplying the features of NFP and GAP. The model architectures for each backbone and pooling variant are trained and evaluated as described above. Results are summarized in Tab. 1.
@@ 4.3 Overall Comparison
Table 1 compares the proposed NFP against global average pooling (GAP) and recent texture pooling methods (Lacunarity [15], Fractal [7], RADAM [22], and DeepTEN [31]) across five datasets and three backbones. Overall, NFP consistently matches or outperforms the strongest baselines while introducing negligible parameter cost. GAP remains competitive on simpler datasets such as PlantVillage and UC Merced, where nearly all methods achieve accuracies above 98–99%. However, it struggles on texture-heavy domains like GTOS-Mobile, particularly with ViT-Tiny (68.64%), since global averaging discards local structural detail. Texture-aware pooling methods, including NFP, address this limitation by explicitly encoding neighborhood-level relationships.
Handcrafted descriptors such as Lacunarity and Fractal pooling provide modest or inconsistent gains over GAP but are limited by fixed statistical formulations. On ResNet18 with GTOS-Mobile, for instance, Lacunarity improves GAP by just over one percentage point (73.12% vs. 72.01%) but remains below NFP (74.52%). On MobileNetV3, Lacunarity actually underperforms GAP on GTOS-Mobile (73.81% vs. 76.27%) and again falls short compared to NFP (76.49%). These results highlight the inconsistent benefits of handcrafted pooling compared to learnable similarity-based features.
RADAM, which relies on stochastic aggregation of activation maps, can outperform NFP in isolated cases. For example, with MobileNetV3 on UC Merced, RADAM slightly exceeds NFP (98.91% vs. 98.49%). However, this comes at the cost of stability: on GTOS-Mobile, NFP is more reliable, surpassing RADAM by +0.6% (76.49% vs. 75.89%). DeepTEN achieves strong performance on large-scale datasets such as RESISC45, reaching 94.95% with MobileNetV3. Yet these gains come with significant parameter overhead (+1.46M compared to GAP). NFP attains nearly the same accuracy (94.80%) with only +0.09M additional parameters, making it a far more efficient option. On ViT-Tiny with GTOS-Mobile, NFP also surpasses DeepTEN by +5.6% (75.47% vs. 69.84%) despite being significantly lighter. Another advantage of NFP is its robustness. Unlike stochastic approaches such as RADAM, which exhibit high variance across runs (e.g., MobileNetV3 on GTOS-Mobile: $\pm 3.58$), NFP produces stable results with consistently low deviation. This reliability, combined with its lightweight design, makes NFP particularly attractive for real-world deployments where memory footprint and stability are critical.
Backbone-specific trends. For CNN-based backbones, NFP provides consistent improvements across datasets, particularly on GTOS-Mobile where texture cues dominate classification. On ResNet18, NFP achieves 74.52%, outperforming Lacunarity (73.12%) and RADAM (69.37%), and shows the highest overall accuracy on UC Merced (98.86%). With MobileNetV3, NFP again delivers the top result on GTOS-Mobile (76.49%), exceeding GAP (76.27%) and clearly surpassing Lacunarity (73.81%). For ViT-Tiny, improvements are smaller on simpler datasets such as PlantVillage and UC Merced but substantial on GTOS-Mobile, where NFP improves over GAP by +6.83%. These results suggest that convolutional backbones benefit most from the neighborhood similarity mechanism, while transformers gain selectively on texture-dominated datasets.
Similarity Measure Accuracy $\pm$ Std (%)
Cosine 98.86 $\pm$ 0.58
Dot Product 98.25 $\pm$ 0.62
Root Mean Square Error (RMSE) 98.02 $\pm$ 0.56
Sharpened Cosine (SCS) 98.02 $\pm$ 0.79
L1 Norm 98.41 $\pm$ 0.40
L2 Norm 98.25 $\pm$ 0.68
Geman-McClure 97.78 $\pm$ 0.56
Scaled Dot Product 98.10 $\pm$ 0.51
Earth Mover’s Distance (EMD) 98.49 $\pm$ 0.59
Canberra 98.17 $\pm$ 0.59
Hellinger 98.25 $\pm$ 0.56
Chi-Squared Type 1 98.10 $\pm$ 0.34
Chi-Squared Type 2 98.02 $\pm$ 0.49
GFC 98.25 $\pm$ 0.68
Pearson Correlation 97.86 $\pm$ 0.51
Jeffrey Divergence 98.13 $\pm$ 0.58
Squared Chord 98.10 $\pm$ 0.89
Smith Similarity 98.02 $\pm$ 0.98
Table 2: Comparison of similarity measures for NFP using ResNet-18 [27] on the UC Merced dataset [30]. Best average result in bold.
(a) Input Image
(b) Average Feature Map
(c) Average NFP
(d) Average Lacunarity
Figure 3: Visualization of feature representations from the first layer of MobileNetV3 [10] on the UC Merced dataset [30], class “intersection.” (b) The standard feature map highlights low-level edges and textures, (c) NFP (cosine similarity) enhances local neighborhood structures, and (d) lacunarity pooling emphasizes spatial gaps and texture distributions. Feature map visualizations are channel-averaged and normalized to the [0,1] range for comparison.
4.4 Similarity Measures

### v3

@@ Abstract
In this work, we propose neighborhood feature pooling (NFP) as a novel texture feature extraction method for remote sensing image classification. The NFP layer captures relationships between neighboring inputs and efficiently aggregates local similarities across feature dimensions. This new approach is implemented using convolutional layers and can be seamlessly integrated into any network. Results comparing the baseline models and the NFP method indicate the potential of this new approach for classification. NFP consistently improves performance across diverse datasets and architectures while maintaining minimal parameter overhead. The code for this work is publicly available [^1] https://github.com/Advanced-Vision-and-Learning-Lab/Neighbour_Feature_Pooling.
@@ 1 Introduction
However, standard CNN architectures can inadvertently lose these cues when using traditional pooling operations (e.g., average and max pooling) [20]. These layers are effective in summarizing information at reduced spatial dimensions, yet typically fail to encode important structural details [32]. This limitation stems from the lack of explicit weighting on neighboring relationships within the feature space. To address this gap, neighborhood feature pooling (NFP) is introduced in this work. This is a novel layer that measures how similar each neighbor is to a center pixel and/or feature vector. NFP can work alongside the usual global average pooling (GAP), adding a texture-aware branch for improved feature representation.
@@ 2 Related Work
(b) An input patch from the UC Merced dataset (“buildings” class) is processed with eight directional difference kernels, which measure relative feature differences between the center pixel and its neighbors (scaled dot product similarity). For visualization, the outputs are normalized to the range [0,1] and computed with a dilation factor of 15, yielding NFP maps that form a multi-channel similarity representation of local texture relationships, later aggregated into the NFP feature vector.
Figure 1: Overview of the NFP framework. (a) NFP example of feature aggregation process. (b) Example with $r=1$, showing the input patch, directional kernels, and the resulting similarity maps that encode local texture relationships. Together, these demonstrate how NFP captures local structural information that is later integrated with global features (Fig. 2).
@@
Texture descriptors in images may be statistical, summarizing pixel-value distributions, or structural, modeling relationships among neighboring pixels [20]. Two structural approaches exploit similarity maps to enrich convolutional features. The Local Similarity Pattern (LSP) layer appends cosine-similarity maps between each pixel and its eight-neighbor window, supplying additional structural cues for stereo-matching networks [13]. In multi-sensor anomalous change detection, a Neighborhood Similarity Feature Space (NSFS) concatenates dot-product similarity responses as extra channels, boosting texture discrimination without resampling [7]. While these methods demonstrate the benefit of neighbor-wise similarity, LSP and NSFS only used cosine similarity and did not investigate the use of other measures to quantify neighborhood relationships as well as these methods were not implemented using convolutional layers that can be readily added to any network in different locations.
@@ 3 Method
Figure 2: Full architecture illustration of the proposed model using MobileNetV3 [9] with NFP. Each block represents a stage in the feature extraction pipeline. After feature maps are extracted from the input, the features are aggregated through two branches: global average pooling (GAP) and NFP. The NFP branch first extracts the similarity maps then the similarity values are aggregated through GAP. The average NFP features are then upsampled to the same dimension using a $1\times 1$ convolution. The final step is for the GAP and NFP feature vectors to be multiplied before being passed into the output classification layer. The $\odot$ symbol denotes element-wise multiplication used to fuse the GAP and NFP feature vectors.
@@ 3.1 Neighborhood Similarity Computation
where $d(\cdot,\cdot)$ is a similarity function such as cosine or dot product. This operation yields a set of $(2r+1)^{2}-1$ scalar values per center pixel, producing similarity feature maps that encode local texture structure, as illustrated in Fig. 1. NFP also employs a predefined similarity function (e.g.,cosine similarity), but applies it locally within each feature map and aggregates the resulting neighborhood responses as a lightweight layer. Since the backbone and projection layers are trained end to end, the feature embeddings entering the similarity computation adapt during learning, so the similarity patterns themselves evolve even though the metric form is fixed. This localized, similarity-driven pooling preserves structural context while adding little overhead, making NFP a practical alternative to methods that require more parameters or codebooks such as DeepTEN and RADAM in remote-sensing classification.
Similarity measures such as cosine or Pearson correlation can yield negative values, which are retained to represent dissimilar or anti-correlated relationships between neighbors. For distance-based metrics such as $L_{p}$ norm, the computed distances are negative so that higher similarity consistently corresponds to larger values (i.e., values closer to zero indicate more similarity). This convention provides a unified interpretation across metrics and ensures that all similarity and distance functions are handled within a common framework. The current NFP layer uses a radius of $r=1$ to capture strictly local relationships while avoiding the quadratic growth in similarity computations and redundancy with receptive field expansion already provided by deeper convolutional layers which results in 8 similarity feature maps given any input feature dimension $d$. Additionally, this feature is inspired from LBP which defines a radius and number of equally spaced pixels, $P$, to define the neighborhood [19]. LBP generally uses a circular neigbhorhood, but the proposed NFP leverages convolutional layers in Pytorch that use a square neighborhood. To keep our method comparable to the original LBP method, a radius of 1 and $P$ = 8 is used to be similar to a $3\times 3$ square neighborhood. The formulation in Eq. 1 thus provides an explicit mechanism for encoding relative feature-space relationships across local neighborhoods, preserving structural texture cues often discarded by global pooling.
@@ 3.2 Similarity Function Options
The similarity function $d(\cdot,\cdot)$ in Eq. 1 can be drawn from a wide set of metrics, each emphasizing different aspects of feature space geometry. The choice of metric can influence NFP by defining how local neighborhood relationships are quantified. We follow the taxonomy of the first three categories defined by Deborah et al. [4] as the feature vectors are similar to hyperspectral signatures used in remote sensing. We also introduced additional similarity metrics inspired by deep learning methods:
@@
Backbone Method RESISC45 UC Merced EuroSAT
ResNet18 GAP 91.28 ± 0.04 (11.20) 96.62 ± 0.19 (11.19) 98.27 ± 0.19 (11.21)
Lacunarity 93.10 ± 0.01 (11.20) 97.30 ± 0.11 (11.19) 98.23 ± 0.33 (11.21)
Fractal 92.29 ± 0.16 (11.46) 97.62 ± 0.34 (11.45) 98.49 ± 0.20 (11.48)
RADAM 92.08 ± 0.15 (11.20) 97.46 ± 0.22 (11.19) 98.40 ± 0.23 (11.21)
DeepTEN 93.19 ± 0.05 (11.96) 97.14 ± 0.58 (11.57) 98.50 ± 0.28 (11.42)
NFP (Ours) 93.22 ± 0.25 (11.20) 98.86 ± 0.58 (11.19) 98.52 ± 0.20 (11.22)
MobileNetV3 GAP 93.48 ± 0.37 (4.25) 97.57 ± 0.57 (4.22) 98.23 ± 0.13 (4.21)
Lacunarity 92.93 ± 0.41 (4.25) 97.94 ± 0.59 (4.22) 98.52 ± 0.32 (4.21)
Fractal 93.55 ± 0.19 (5.17) 97.94 ± 0.22 (5.15) 98.35 ± 0.15 (5.14)
RADAM 85.59 ± 2.61 (4.25) 97.78 ± 1.81 (4.22) 98.27 ± 0.05 (4.21)
DeepTEN 94.95 ± 0.09 (5.68) 98.49 ± 0.31 (4.94) 98.35 ± 0.23 (4.60)
NFP (Ours) 94.80 ± 0.21 (4.25) 98.49 ± 0.28 (4.23) 98.53 ± 0.12 (4.22)
ConvNeXt-Atto GAP 93.90 ± 0.27 (3.39) 96.19 ± 0.78 (3.38) 96.33 ± 0.26 (3.38)
Lacunarity 93.25 ± 0.36 (3.39) 94.37 ± 0.66 (3.38) 96.90 ± 0.36 (3.38)
Fractal 90.58 ± 0.56 (3.43) 95.79 ± 0.22 (3.42) 96.38 ± 0.35 (3.42)
RADAM 91.61 ± 0.84 (3.39) 95.63 ± 0.22 (3.38) 97.16 ± 0.26 (3.38)
DeepTEN 94.27 ± 0.31 (3.85) 96.27 ± 0.30 (3.60) 97.76 ± 0.17 (3.49)
NFP (Ours) 94.71 ± 0.21 (3.39) 97.30 ± 1.07 (3.38) 97.65 ± 0.26 (3.38)
Table 1: Classification accuracy (%) and model size (in millions of parameters) across three remote-sensing datasets. Each cell shows average accuracy $\pm$ 1 standard deviation, followed by parameter count in parentheses. For each backbone architecture, the best performing (i.e., highest average accuracy) method on each dataset is bolded.
@@ 3.3 Model Architecture Integration
The affinity stack is spatially averaged and projected to a vector in $\mathbb{R}^{C^{\prime}}$ with $C^{\prime}=C$. A $1\times 1$ convolution is applied only if $C^{\prime}\neq C$. The resulting vector is fused with the backbone’s GAP representation via element-wise multiplication. Because both streams are pooled to vectors, the difference between $(H,W)$ and $(H^{\prime},W^{\prime})$ does not affect the fusion. This design preserves pretrained backbone weights and introduces only a lightweight projection layer as additional learnable overhead, making NFP broadly applicable across different architectures.
Figure 3: t-SNE visualizations of penultimate-layer features extracted from models trained on RESISC45 (45 scene classes) using different pooling methods: GAP, NFP (Ours), DeepTEN, Fractal, Lacunarity, and RADAM. Each point represents a test image, colored by its ground-truth class label. The Silhouette Score, computed from the original feature space, is reported beneath each panel and reflects the compactness and separability of the classes. Among the compared approaches, NFP achieves the highest Silhouette Score, indicating its superior ability to generate more discriminative and well-separated feature embeddings for remote sensing scene classification. All visualizations were generated using a shared random seed to ensure a fair and consistent comparison across methods.
4 Experimental Results and Discussion
@@ 4.1 Experimental setup
Three public remote sensing datasets are used for evaluation: UC Merced Land Use [29], RESISC45 [3], and EuroSAT [8]. All experiments follow the standard train/validation/test protocols described by Neumann et al. [17]. To ensure fairness across baselines, all models are trained using identical data splits and preprocessing pipelines. Since EuroSAT contains 13 spectral bands, the ImageNet-initialized first layer is adapted using mean channel-wise initialization, which shows the most consistent performance among tested strategies in Sec. 4.7.
@@ 4.2 Backbone Networks
The NFP layer is evaluated across three convolutional backbone architectures: ConvNeXt-Atto [14], MobileNetV3-Large [9], and ResNet-18 [26]. All models are initialized with pre-trained weights from the PyTorch Image Models (timm) library [27] and fine-tuned on each dataset. The NFP module is integrated into each backbone according to the scheme described in Sec. 3.3. In all cases, NFP is placed after the final convolutional feature extraction stage, followed by a linear classification head. A $1{\times}1$ convolution is applied after NFP to align the number of feature maps to those produced by GAP. Classification is performed by element-wise multiplication of the NFP and GAP feature vectors, followed by a linear classifier. The model architectures for each backbone and pooling variant are trained and evaluated as described above, and results are summarized in Tab. 1.
@@ 4.3 Overall Comparison
Table 1 compares the proposed NFP against GAP and recent texture pooling methods (Lacunarity [15], Fractal [5], RADAM [23], and DeepTEN [30]) across three remote-sensing datasets and three backbones (ResNet18, MobileNetV3-Large, and ConvNeXt-Atto). Although NFP is conceptually related to neighborhood-similarity modules such as LSP and NSFS and also relies on a cosine-based similarity operator, those modules were originally designed for stereo matching and bi-temporal change detection and are tightly coupled to task-specific architectures. Adapting them into lightweight pooling heads for single-image scene classification would require substantial redesign, so here we focus on controlled comparisons against established texture pooling heads that can be used as drop-in replacements on standard backbones. Accordingly, NFP is best viewed as a lightweight pooling reformulation of neighborhood similarity rather than a new backbone or feature extraction paradigm.
Across all backbone–dataset combinations, NFP is competitive with the strongest baseline and achieves the best or statistically tied accuracy in most cases, while keeping the parameter count essentially identical to the GAP baseline. In the remaining settings, NFP trails the top method by only a very small margin. On simpler benchmarks such as UC Merced, GAP remains a strong baseline and nearly all methods approach saturated performance, but texture-aware pooling still offers a more explicit modeling of neighborhood-level relationships that are attenuated by global average pooling. Methods inspired by handcrafted descriptors such as Lacunarity and Fractal pooling provide modest or inconsistent gains and are constrained by fixed statistical formulations, whereas NFP and other learnable similarity-based pooling heads adapt to the data and yield more reliable improvements across settings.
RADAM, which aggregates activation maps stochastically, can outperform NFP in isolated configurations but exhibits noticeably higher variance across runs, especially on the more challenging dataset–backbone pairs. DeepTEN attains slightly higher accuracy than NFP in a few cases, but these gains are confined to sub-percent differences and come at a substantially higher parameter cost, since DeepTEN consistently enlarges the model relative to the baseline, whereas NFP preserves the lightweight footprint of GAP. This trade-off highlights NFP as an attractive middle ground that captures texture-aware information with minimal complexity and stable optimization behaviour.
Backbone-specific trends For ResNet18, NFP improves over GAP and handcrafted pooling on all three datasets without increasing model size. With MobileNetV3, NFP matches the strongest texture baseline on the easiest dataset and remains very close to the best method on the others, again with negligible parameter overhead. For ConvNeXt-Atto, NFP yields the best performance on two of the three datasets and remains within a narrow margin of the top method on the remaining one, offering a favorable balance between accuracy and efficiency for compact CNN-style backbones.
t-SNE Visualization Analysis Penultimate-layer features from each model are visualized using t-SNE [24] on the RESISC45 dataset. Figure 3 presents two-dimensional projections for GAP, Fractal, Lacunarity, RADAM, DeepTEN, and the proposed NFP, with each point colored according to its ground-truth scene class. All visualizations are generated using identical t-SNE hyperparameters and a shared random seed to ensure a fair comparison. The corresponding Silhouette Scores [22] are reported beneath each plot to quantify class compactness and separability. NFP achieves the highest score (0.5600), followed by GAP (0.5160) and DeepTEN (0.5226). In contrast, Lacunarity (0.4343), RADAM (0.3350), and Fractal (0.3136) exhibit significantly lower values, indicating less class separability and compactness. Both the qualitative t-SNE projections and the quantitative Silhouette Scores demonstrate that NFP produces more compact and better-separated feature embeddings for the RESISC45 dataset.
Method FLOPs Memory Latency
GAP 75.80 331 $5.42\pm 0.83$
Lacunarity 75.80 549 $4.89\pm 0.01$
Fractal 76.34 593 $5.21\pm 0.13$
RADAM 75.80 636 $31.94\pm 0.11$
DeepTEN 75.81 688 $6.20\pm 0.01$
NFP 75.89 375 $5.36\pm 0.01$
Table 2: Efficiency comparison of pooling methods on ResNet18 (input size $\times256$, batch size 32). FLOPs are reported per forward pass (batch size 32) in gigaflops (G), memory as peak graphics processing unit (GPU) allocation in megabytes (MB), and latency as average per-batch inference time in milliseconds (ms) over multiple forward passes on a single GPU. NFP uses cosine similarity.
Figure 4: Confusion matrices on the RESISC45 dataset using a ConvNeXt-Atto backbone with three different pooling strategies: GAP, DeepTEN, and the proposed NFP. The NFP variant demonstrates the strongest diagonal dominance and the lowest degree of inter-class confusion.
Computational efficiency Computational cost is quantified in terms of floating-point operations (FLOPs), peak graphics processing unit (GPU) memory usage, and inference latency for each pooling head combined with a ResNet18 backbone on $\times256$ inputs and a batch size of 32. As reported in Tab. 2, replacing GAP with NFP (using cosine similarity) yields almost identical FLOPs and latency, with only a modest increase in memory usage. In contrast, Lacunarity, Fractal, DeepTEN, and RADAM exhibit noticeably higher memory requirements, with RADAM in particular showing increased latency. Together with the accuracy gains over GAP in Tab. 1, these measurements indicate that NFP improves classification performance while largely preserving the computational characteristics of standard GAP.
Confusion Matrix Analysis Figure 4 compares the confusion matrices of ConvNeXt backbone using GAP (baseline), DeepTEN, and the proposed NFP on the RESISC45 dataset. The baseline model shows noticeable off-diagonal activations, indicating confusion between visually similar scene categories such as river and lake, terrace and mountain, and dense residential and commercial areas. DeepTEN partially reduces this confusion; however, residual misclassifications remain in several texture-rich and structurally complex classes.
In contrast, the NFP model shows a much cleaner diagonal structure with substantially reduced inter-class confusion. This improvement is most evident in categories with complex spatial layouts and fine-grained textures, including island, river, mobile home park, snowberg, terrace, and wetland. These results highlight the primary advantage of NFP: its ability to capture neighborhood-level similarities and preserve local spatial relationships, leading to stronger class separability and more reliable scene discrimination than both GAP and DeepTEN.
4.4 Ablation: Similarity Measures